OpenStudy (anonymous):

I have a question about similare shapes. The question gives me two triangles and the information of the triangles as follows: Triangle A: Base length of 20cm Area of 187.5cm^2 Triangle B: Base length of 30cm The question asks me to find the area of triangle B. I will post what I tried to do as a comment :)

OpenStudy (anonymous):

What I tried: \[\frac{ 30^{2} }{ 20^{2} }= \frac{ x }{ 187.5 }\] \[\frac{ 900 }{ 400 } x 187.5= 421.875\]

OpenStudy (anonymous):

That last equation should be a multiplication sign not an x (oops sorry)

OpenStudy (anonymous):

You don't need to square 20 & 30. The first thing you should do is find the height of the first triangle using the area formula A=1/2 *base*height 187.5=1/2*20*x x=18.75 Now, you'd use proportions 20/18.75=30/x Solve for x

OpenStudy (anonymous):

I get 32?

OpenStudy (md152727):

x=28.124991

OpenStudy (anonymous):

In my text book the answer is supposed to be 125cm^2

OpenStudy (anonymous):

I can post a picture of the question if that would help? Its possible I didnt explain things very well

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

Ok, coming up

OpenStudy (anonymous):

It is question a)

OpenStudy (anonymous):

This is an example given and how i tried to solve it

OpenStudy (anonymous):

You explained it right. We still get x=28.125 though, which would make the are of triangle B 421.875cm^2 Sorry we couldn't help.

OpenStudy (anonymous):

Thats ok, thank you for trying :)